

********************************************************************
UARS WEEKLY STATUS REPORT - 09 APR 1999
********************************************************************
***********
SOLSTICE
***********
SOLSTICE began day only operations on Sunday, April 4 (99094) and will
return to normal operations at the begining of day 99101 (Sunday, April 11).
Full solar data sets are retrieved during this mode, however no stellar
calibration obervations are possible.
***********
SUSIM
***********
SUSIM turned off on April 3. It is scheduled to turn on April 11.
Reprocessing of the SUSIM data to version 20 continues.
***********
PEM
***********
PEM/AXIS was deactivated on April 3. While AXIS was activated, measured
X-rays were inverted to determine the incident electron population and
this electron population was used to determine the amount of energy
deposited in the atmosphere by electrons and bremsstrahlung X-rays.
PEM is active under PMON control. Energy processor #6 for the
HEPS-1 unit has been found to not activate correctly. Data affected
are from the upper energy EE channel for Head-2 of HEPS-1 (that is the
+15 deg electrons above 300 keV). Tests are being conducted on this
energy processor and other energy processors to determine conflicts.
Measurements from this sensor will be activated as soon as possible.
The PEM team and FOT are convinced that the entanglement of energy
processor #6 is related to OBC control and not a failure of that
energy processor. Early attempts at adjusting OBC command timing
have not been successful. The FOT and the PEM team is attempting to
refine their diagnostics and confine the problem.
While under PMON control, the PEM instrument computes energy deposition
from the particle spectrometers.
***********
HALOE
***********
HALOE continues successful operations with all engineering parameters
within nominal operating ranges. Science data products are of excellent
quality.
***********
MLS
***********
MLS was reactivated successfully on Apr 4 and is scheduled to operate through
Apr 10.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
WINDII is currently off, but expecting to be activated for
the period April 13 to 19. WINDII reprocessing is up to Day 233.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
1 day HAL (4.81) PROD (L1-L3) (1R-1C=0L)
1 day PEM (3.32) PROD (L1-L3) (1R-1C=0L)
50 days PEM (3.32) PROD (OA-L2) (760R-428C-1F=331L)
1 day SOL (4.62) TEST (L2-L3 - DEBUG) (2R-1C-1F=0L)
1 day SOL (4.62) PROD (L1-L3) (1R-1C=0L)
49 days SUS (7.1) PROD (L1 - SUBSET) (49R-49C=0L)
1 day SUS (7.1) PROD (L3) (1R-1C=0L)
253 days WIN (5.11) PROD (LOG) (2014R-669C=1345L)
75 days WIN (5.11) PROD (L1-L3) (2014R-149C=1865L)
UARS CDHF REHOST:
We are rehosting the current UARS CDHF cluster of VAX and ALPHA machines to all
ALPHA's. The new CDHF consists of two ALPHA 4000s and a Tape Library with 6TB
of storage. We plan on transitioning to the 4000s for current day processing
in September 1999. Reprocessing on the 4000s is dependent on how quickly we
can migrate off the current storage library. We shall keep you informed of the
transition plan and timeline. This effort will increase processing power,
decrease footprint, and decrease cost.
The CDHF will load February 1999 current day DAAC data in April 1999.
The CDHF set up and executed a SOLSTICE day that used excessive wallclock
time with an Executable supplied by the UCSS Developer. Results are
being evaluated by developer.
CDHF pending software upgrades: DEC AXP Fortran 6.5, DEC AXP OPEN
VMS 6.2, OpenIngres.
***********
FOT
***********
This report covers 27 March 1999 (Orbit 41200, GMT Day 086 through
02 April, GMT Day 092).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing) and
is performing nominally under two-battery operations. The instrument
operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM 089/0247 41247 AXIS On
HALOE 090/2233 41274 On
The beta angle ranged between the angles of 7.1 degrees to 19.8 degrees for
this report period. The beta angle is now increasing toward a maximum
angle of 77 degrees on 21 May 1999 (DOY 141).
A Forward-to-Reverse Yaw Around maneuver (number 77) was successfully
executed in Orbit 41211 beginning 086/15:55:48 and completing at
086/16:32:58 on 27 March 1999. The SSPP Instruments (ACRIM, SOLSTICE,
SUSIM) were enabled as well as HRDI and PEM in day only operations. Due to
the SA position at 269 degrees, the effective night length of the
spacecraft after the maneuver was approximately 50 minutes. Post maneuver
analysis of data indicate that during the Yaw Around, the end-of-night load
bus voltage was 24.5 volts and the average depths of discharge (DODs) on
Batteries 2 and 3 were approximately 22.0 and 22.5 percent respectively.
The next (reverse-to-forward) YAW is scheduled for 30 April 1999.
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage.
Battery temperatures are stable with a temperature delta between Battery 2
and Battery 3 of 0.9 to 1.3 degrees C. Additional battery performance data
for each battery on 26 March 1999 and 02 April 1999 are provided in
Appendix D.
The clock error ranged from -10.8 to 10.1 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT for
this report. A summary is listed below and the details of ALERT and Out-
Of-Limits occurrences are listed in Appendix B.
% ALERTS = 01
% OUT-OF-LIMITS = 04
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
% FOT 03
% HALOE 02
% PEM 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
% FOT 00
Two (2) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
Transponder A: 99/054 - 2287.496948
Transponder B: 99/055 - 2287.498731
GROUND SYSTEM OPERATION
No (0) operational support problems occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from
mission start through UTC Day 99/091 show the observatory data loss to be
22 hours, 46 minutes, 37.396 seconds (an increase of 1.696 seconds since
last report period). This is a 0.0345 percent data loss which equals a
99.9655 percent data capture for the mission.
The increase of 1.696 seconds occurred over three (3) days:
7 1.024 seconds on Day 085 due to data hit,
7 0.448 seconds on Day 086 due to data hit, and
7 0.224 seconds on Day 088 due to data hit.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT continues to support the PEM team with diagnosis of the Energy
Processor 6 status error.
The FOT continues to investigate the feasibility of a new PMON action to
change battery V/T curves based on loading to mitigate battery risks in sun
pointing safehold.
YEAR 2000 OPERABILITY SUPPORT - No change
UARS PROPULSION MODULE B/D THRUSTER LINE FILL
Planning continues for the simulator testing.
MEETINGS
FOT supported an IMOC meeting on 31 March 1999.
FOT supported a Leonid shower (second event) meeting conducted on 01 April
1999 by Omitron personnel. The purpose of the meeting was to provide
answers to a set of questions being used as input to a composite report of
previous experience, future modifications, additions, etc.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw Around Maneuver (reverse-to-forward) is scheduled for 30
April 1999 (DOY 120).
PERSONNEL CHANGES
None.
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 02 April 1999
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
41211 086/1708 99-015 MACS YAW MANEUVER
41247 089/0247 99-116 PEM AXIS Power ON
41274 090/2233 99-118 HALOE Power ON
41279 091/0705 99-118 HALOE Power ON - Full Power
41280 091/0849 99-119 OBC Clk Adj fr 0.00 to 20.6 msec/day
41298 092/1230 99-120 OBC Clk Adj fr 20.6 to 0.00 msec/day
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 02 April 1999
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
HRPSEQBIMON Y-HI PSE QB Current 41249
HAH28VA R-LO HALOE Htr Mtr V 41275-9
HRPSEQBIMON Y-HI PSE QB I Mon 41279-80
NBRECBI Y-HI NBTR B Current 41280
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
SSSWLIMFAIL SSPP Software Limit Fail 41216
_______________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 02 April 1999
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
41280 091/08:56 (01 Apr 99) from 00.0 to 20.6 msec/day
41298 092/12:32 (02 Apr 99) from 20.6 to 00.0 msec/day
_______________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 02 April 1999
26 Mar 1999 - GMT Day 085 Beta = 7.1 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 -33.6/-50.4 -3.68/-4.94 +0.0/ +0.0 16.8 0.000 0.0
2 +44.8/-11.2 +4.83/+3.55 +30.4/-13.6 25.6 1.016 20.6
3 +89.6/0.0 +3.87/+2.60 +30.4/-14.4 25.4 1.014 20.4
02 Apr 1999 - GMT Day 092 Beta = 19.8 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 -33.6/-56.0 -3.05/-4.31 +0.0/ +0.0 16.6 0.000 0.0
2 +89.6/ 0.0 +6.11/+5.15 +30.8/-15.2 24.8 1.024 21.4
3 +56.0/ -5.6 +5.15/+3.87 +30.4/-16.0 24.5 1.023 21.5
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes V/T 5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 02 April 1999
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
None this report period
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
None this report period
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------- ------ ----- ---- --- -----------------
None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-011 FOT see list Mar 99 14170 Generic Late Acq
99-014 FOT 41189 085/0548 N/A CRC Errors in PBK
SEVNTFW99/14